a cargo ship with a mass of 5×105 kg floating in water carrying loads with unknown weight, WL. Suppose the cargo ship displaces an amount of 6300 m3 seawater while carrying the loads, determine: (a) the buoyant force exerted on the cargo ship. (b) the maximum weight of loads, WL the cargo ship can carry before it sinks


1
Expert's answer
2021-09-22T07:05:53-0400

(a) The buoyant force exerted on the cargo ship is


"F_b=\rho gV=1000·9.8·6300=61740000\text{ N}."

(b) The maximum weight of loads, WL, the cargo ship can carry before it sinks, is


"W_L=F_b-mg=\\=61740000-5·10^5·9.8=56840000\text{ N}."